AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 454     Entry time: Mon Oct 16 08:28:02 2023
Author: CB 
Subject: NEG activation -1 

9:00 - Michael's online vacuum pressure plot is no longer working - an old image is being displayed. CARME temperatures around 80-85C. Lowered heater temperatures to 150, 100, 150 C for safety. Lowered jacket temperature to 80C

9:15 - Flashed ion pumps several times. Small pressure increase, up to ~7E-9 but not macroscopic

9:30 - Started D-2000 #1,2. T=400, tramp 30, thold 90 (this was a mistake 1,3 are behind the detectors, 2,4 are to the side. They should ideally be activated together. Got confused. Not a big deal.)

9:50 - Online vacuum plot is working again. Pressure changing, internals around 80C

10:00 - Internal temperature still stable at 80C. Started D-2000 #3,4. T=400, tramp 30, thold 90

11:30 - For unclerar reasons the pressure has been increasing over the last half hour, after the four NEGs were activated. Nikos and the alignment group were adjusting the feet of the interaction chamber to improve the alignment. Hopefully this is not a leak. Section 10 was apparently vented overnight, and may be in air at the moment.

11:45 - D2000 activation complete. Set temperature to 200C for all four of them. Altough the temperature is still >375C, the pressure drops very quickly. This is good, but I have not observed this behaviour before.

13:10 - Pressure and temperature mostly stable. Turning on UHV1400 groups #1, 2 with at P=250W, tramp 30s, thold 90s.

13:35 - No effect on DSSD temperature. Increasing P->300W as per previous activation procedure.

13:43 - Still no sizeable effect. Increased P->350 W

15:00 - Completed activation of UHV1400 sets #1,2. Set power to 100W. Activated sets #3,4 at P=250W, same settings as other two.

15:30 - Increased UHV1400 #3,4 to P=300 W. Temperatures OK.

15:37 - Increased UHV1400 #3,4 to P=350 W. Temperatures increasing but still acceptable.

16:06 - Temperatures are starting to get a bit too high. Decreased P->300W. Also decreased heaters to 125, 80, 125C and baking jacket to 70 C.

16:26 - Decreased UHV1400 #3,4 to P=250W.

17:05 - Completed activation of UHV1400 sets #3,4. Set power to 100W. Planning to leave stabilising overnight.

Attachment 1: Screenshot_2023-10-16_17-17-30.png  211 kB  Uploaded Mon Oct 16 17:17:38 2023  | Hide | Hide all
Screenshot_2023-10-16_17-17-30.png
ELOG V3.1.4-unknown